OpenStudy (anonymous):

Find the differential equation of all tangent lines of the parabola y=x^2

OpenStudy (anonymous):

Find the derivative with respect to x.

OpenStudy (jhannybean):

\[\large \frac{dy}{dx}(y=x^2)\]\[\large \frac{d}{dx}(y)=\frac{d}{dx}(x^2)\]\[\large y'= 2x\]